### v1

# FastVID: Dynamic Density Pruning for Fast Video Large Language Models

Abstract
Video Large Language Models have shown impressive capabilities in video comprehension, yet their practical deployment is hindered by substantial inference costs caused by redundant video tokens. Existing pruning techniques fail to fully exploit the spatiotemporal redundancy inherent in video data. To bridge this gap, we perform a systematic analysis of video redundancy from two perspectives: temporal context and visual context. Leveraging this insight, we propose Dynamic Density Pruning for Fast Video LLMs termed FastVID. Specifically, FastVID dynamically partitions videos into temporally ordered segments to preserve temporal structure and applies a density-based token pruning strategy to maintain essential visual information. Our method significantly reduces computational overhead while maintaining temporal and visual integrity. Extensive evaluations show that FastVID achieves state-of-the-art performance across various short- and long-video benchmarks on leading Video LLMs, including LLaVA-OneVision and LLaVA-Video. Notably, FastVID effectively prunes $\textbf{90}\%$ of video tokens while retaining $\textbf{98.0}\%$ of LLaVA-OneVision’s original performance. The code is available at https://github.com/LunarShen/FastVID.
1 Introduction
Video Large Language Models (Video LLMs) [5, 17, 14, 43, 35] have gained prominence for video comprehension. These models process videos by encoding sampled frames into video tokens through a vision encoder. These tokens are typically pooled before being integrated with query tokens and passed to an LLM for response generation. Despite their effectiveness, the high computational cost of inference remains a significant challenge. While some methods [29, 22, 37, 13, 28, 45] introduce trainable compression modules to mitigate this issue, they require extensive retraining, leading to high training costs. In contrast, we focus on an inference-time acceleration strategy that enhances efficiency without requiring additional training.
Figure 1: Effective token compression in Video LLMs relies on preserving both temporal and visual context. (a) shows the impact of temporal disruptions on video understanding, emphasizing the importance of temporal structure preservation. (b–d) show the effects of different token selection strategies for visual merging (Patches that share the same inner and border color are merged.), indicating that Density-based Sampling retains both distinctive and representative context while effectively reducing redundancy.
This computational burden is primarily caused by the high volume of video tokens, making effective token compression essential. While prior image compression methods [4, 26, 42, 38, 3] reduce redundancy within a single image, they fail to exploit temporal dependencies across frames. Meanwhile, video compression methods like DyCoke [31] and PruneVid [11] primarily prune at the pre-filling and decoding stages, treating video tokens like textual tokens and neglecting video-specific characteristics. There remains a notable gap in the systematic exploration of video redundancy itself. A more video-aware pruning approach is necessary to explicitly exploit spatiotemporal redundancy while preserving essential semantic content.
To address this limitation, we focus on minimizing spatiotemporal redundancy at the input stage. In particular, we systematically analyze video redundancy from two key perspectives: temporal context and visual context.
Temporal context is fundamental to video understanding, as the order and continuity of frames directly influence semantic interpretation. As depicted in Figure 1a, disrupting frame order (shuffled) or omitting frames (incomplete) leads to incorrect comprehension, highlighting the necessity of preserving temporal structure. To achieve this, we segment the video into temporally ordered segments, grouping highly similar consecutive frames. Pruning is applied within each high-redundancy segment but not across segments, preserving the essential temporal structure for accurate video understanding.
Within each segment, it is crucial to merge redundant tokens while preserving visual context. A state-of-the-art method VisionZip [38] applies uniform token sampling followed by merging redundant tokens [3] as shown in Figure 1c. However, this strategy is highly sensitive to token selection, potentially leading to the loss of important details, such as the guitar body being incorrectly merged with the background. To address this, we propose a density-based token sampling in Figure 1d. Specifically, high-density tokens, surrounded by numerous similar tokens, serve as candidates for selection. We select density peak tokens as anchors, which have a higher density than their neighbors and by a relatively large distance from tokens with higher densities [24]. This strategy ensures that the selected tokens are both representative and distinctive, effectively preserving segment visual context while reducing redundancy.
Building on these insights, we propose Dynamic Density Pruning for Fast Video LLMs (FastVID), which optimizes video pruning by dynamically segmenting frames and applying density-based pruning within each segment. We begin with Dynamic Temporal Segmentation, which adaptively partitions videos into segments. Within each segment, we introduce Density Spatiotemporal Pruning to retain both global visual context and salient details. FastVID significantly accelerates Video LLM inference while preserving both temporal and visual integrity.
To assess the generalization capability of our approach, we evaluate it on LLaVA-OneVision [14] and LLaVA-Video [43], two leading Video LLMs. Notably, LLaVA-Video11footnotemark: 1 employs newline tokens based on the positional information of video tokens, which is an aspect that many pruning methods do not explicitly consider. Our method effectively adapts to this design, making it particularly well-suited for LLaVA-Video. To evaluate the effectiveness of our method, we perform extensive experiments on MVBench [16], LongVideoBench [36], MLVU [44], and VideoMME [10]. These benchmarks cover a wide range of video complexities and durations, ensuring a comprehensive evaluation. Notably, FastVID prunes 90% of video tokens while maintaining 98.0% of LLaVA-OneVision’s original performance across all benchmarks.
The main contributions are summarized as follows:
  • We analyze Video LLM compression from both temporal and visual context perspectives, emphasizing the importance of maintaining temporal and visual integrity.
  • We propose FastVID, a novel pruning framework that employs Dynamic Temporal Segmentation to partition videos into temporally ordered segments and Density Spatiotemporal Pruning to retain global segment information and key details.
  • Our FastVID generalizes across different Video LLM architectures and achieves state-of-the-art performance on diverse video benchmarks.
2 Related Work
Video LLMs. With the rapid advancement of LMMs [1, 6, 32] and MLLMs [33, 19, 20, 15, 2, 21], there has been growing interest in Video LLMs. Existing Video LLMs can be broadly categorized based on how they process video tokens: general Video LLMs and Video LLMs with training-time compression.
General Video LLMs [18, 5, 17, 14, 43, 35] directly process raw video tokens or apply pooling. Video-LLaVA [17] leverages shared projection layers to obtain unified visual representations. LLaVA-OneVision [14] demonstrates strong video understanding through task transfer from images. LLaVA-Video [43] creates a high-quality synthetic dataset for video instruction-following. To better capture the spatiotemporal structure of video, some models [43, 35] introduce additional designs for video positional information. Qwen2-VL [35] propose M-RoPE to decompose rotary embedding into temporal, height, and width components. LLaVA-Video [^1] https://github.com/LLaVA-VL/LLaVA-NeXT introduces newline tokens to distinguish spatial and temporal positions effectively.
Video LLMs with training-time compression [29, 22, 37, 13, 28, 45] aim to significantly reduce the number of video tokens, enabling longer video sequences. PLLaVA [37] introduces an adaptive average structure pooling to extend image LLMs. Video-ChatGPT [22] extracts both spatial and temporal features through temporal and spatial pooling respectively. Chat-UniVi [13] progressively clusters visual tokens and provides multi-scale features. MovieChat [29] introduces a memory management mechanism to enhance long video understanding. LongVU [28] employs cross-modal query and inter-frame dependencies to adaptively reduce video redundancy. Apollo [45] explores scaling consistency and uses the Perceiver Resampler [12].
Figure 2: The overview of our FastVID. DySeg dynamically partitions video tokens into temporally ordered, high-redundancy segments, while STPrune performs density-based pruning within each segment, enhancing efficiency without loss of key information.
However, general Video LLMs remain the dominant paradigm, with LLaVA-OneVision being widely adopted due to its adaptability and superior performance. Moreover, Video LLMs with training-time compression often exhibit lower performance compared to their general counterparts. For example, Apollo-7B with trainable compression achieves $61.3$ on VideoMME (wo sub.), whereas LLaVA-Video-7B achieves $63.3$.
Therefore, we focus on inference-time acceleration for general Video LLMs, aiming to improve computational efficiency while maintaining strong performance. To evaluate our method, we select LLaVA-OneVision and LLaVA-Video as representative models.
Token Compression. Token compression has emerged as an effective approach to reduce computational complexity in transformer architectures, such as ViT [8] and CLIP [23]. ToMe [3] progressively merges fixed spatial tokens, while TempMe [27] extends this concept by merging neighboring clips to minimize temporal redundancy.
Recent studies [4, 26, 42, 38, 3] primarily focus on spatial token reduction to accelerate Image LLMs. FastV [4] selects text-relevant tokens at shallow layers of the LLM. LLaVA-PruMerge [26] uses attention scores from the [CLS] token to prune spatial redundancy. SparseVLM [42] proposes a token recycling strategy to aggregate and reconstruct tokens to be pruned. VisionZip [38] reduces visual redundancy in the vision encoders. However, these methods overlook the temporal relationships across frames.
Due to the high volume of video tokens in Video LLMs, recent video compression methods [31, 11] have gained increasing attention. DyCoke [31] merges tokens across frames and applies dynamic KV cache reduction. PruneVID [11] clusters video tokens and selects those most relevant to query tokens. However, pruning in these methods predominantly occurs during the prefilling and decoding stages, where spatiotemporal characteristics of video data are not adequately considered.
In this paper, we focus on temporal and visual context for spatiotemporal token pruning at the input stage. By reducing video tokens before LLM processing, our FastVID significantly enhances computational efficiency while facilitating easy deployment, including compatibility with FlashAttention [7], KV cache, and a plug-and-play design for seamless integration into existing Video LLMs.
3 Methodology
3.1 FastVID
Video LLMs face significant inference challenges due to the large number of frames in a video and the numerous vision tokens within each frame. Typically, the total token count often exceeds the LLM’s context length, making direct processing computationally infeasible. Although uniform frame sampling and token pooling methods are commonly employed to reduce complexity, they still leave a substantial number of redundant video tokens (e.g., $32\times 196=6272$ video tokens in LLaVA-OneVision [14]).
To address these limitations, we propose FastVID, a novel method that accelerates inference in Video LLMs by reducing redundant tokens while preserving critical temporal and visual context. First, we introduce Dynamic Temporal Segmentation (DySeg) in Section 3.2, which partitions a video into temporally ordered segments with high redundancy. Within each segment, we introduce Density Spatiotemporal Pruning (STPrune) in Section 3.3, which retains essential visual context and salient details.
Figure 2 presents an overview of FastVID. Given an input video, $F$ frames are uniformly sampled (e.g., 32 in LLaVA-OneVision, 64 in LLaVA-Video). Each frame is individually processed by the vision encoder. The extracted tokens are projected and pooled into a video token sequence. With FastVID, we effectively eliminate redundant tokens while preserving critical information. The retained video sequence, combined with query tokens, is then fed into the LLM to generate the response.
Figure 3: Visualization of different segmentation methods on 12 sampled frames from a video example. Each frame’s top-left corner displays its sequential index. (b) Fixed-interval Segmentation struggles to maintain high intra-segment similarity, leading to visually diverse frames within the same segment. (c) Cluster-based Segmentation disrupts temporal order, grouping frames from different time periods to the same segment. (d) Our DySeg adaptively partitions the video, preserving both temporal structure and high intra-segment similarity. Please refer to Table  2 for a detailed quantitative comparison.
3.2 Dynamic Temporal Segmentation
Section 1 highlights the importance of video segmentation in maintaining temporal order across segments while ensuring high similarity within each segment.
As shown in Figure 3b and 3c, there are two common segmentation approaches for sampled frame sequences. In Figure 3b, Fixed-interval Segmentation partitions the sequence into segments of a fixed length, preserving temporal order but potentially grouping visually dissimilar frames. Figure 3c shows Cluster-based Segmentation, where frames are grouped into three clusters based on frame similarity. However, this approach suffers from a predefined cluster number, leading to ineffective segmentation when video complexity varies. As a result, the first segment contains similar objects but different scenes. Furthermore, clustering may disrupt the temporal order by ignoring critical temporal information, such as omitting key frames (e.g., the ninth frame) and incorrectly grouping frames from different time periods into the first segment.
To address these limitations, we propose Dynamic Temporal Segmentation, a simple yet effective method that adaptively refines segment boundaries according to video complexity. DySeg achieves both temporal structure and high intra-segment similarity, generating fewer partitions for simple scenes and finer ones for more complex scenes.
To maintain high similarity within each segment, we aim to minimize similarity between adjacent segments. Thus, we segment the video based on transition similarity between adjacent frames. Specifically, we utilize global frame features $\mathbf{f}$ to compute the cosine similarity:
$t_{i}=\cos(\mathbf{f}_{i},\mathbf{f}_{i+1}),\quad i=1,\dots,F-1,$
$\mathbf{T}=\{t_{1},t_{2},\dots,t_{F-1}\},$ (1)
where $t_{i}$ denotes the transition similarity between the $i$-th and $(i+1)$-th frame. $\mathbf{T}$ denotes the $F-1$ transition similarities for a sequence of $F$ frames. To achieve fine-grained segmentation, we select transitions that satisfy the following conditions:
$\mathbf{S}_{1}=\operatorname{arg\,min}_{c-1}\mathbf{T},\quad\mathbf{S}_{2}=\{i\mid t_{i}<\tau,\;t_{i}\in\mathbf{T}\},$
$\mathbf{S}=\mathbf{S}_{1}\cup\mathbf{S}_{2},$ (2)
where $c$ denotes the minimum number of segments and $\tau$ denotes the threshold for transition similarity. $\mathbf{S}_{1}$ denotes the $c-1$ least similar frame transitions, while $\mathbf{S}_{2}$ denotes transitions where similarity falls below a fixed threshold $\tau$. The final segmentation is determined by the union $\mathbf{S}$. In simple videos with minimal shot transitions, $\mathbf{S}_{2}$ is often empty. In such cases, $\mathbf{S}_{1}$ enables finer segmentation by distinguishing subtle temporal changes. For complex videos, $\mathbf{S}_{1}$ is typically a subset of $\mathbf{S}_{2}$, where $\mathbf{S}_{2}$ ensures that adjacent frames with similarity below $\tau$ are assigned to different segments. In Figure 3d, our DySeg effectively segments videos by dynamically adjusting granularity in a simple yet effective manner, outperforming Fixed-interval and Clustering-based methods.
3.3 Density Spatiotemporal Pruning
After obtaining segments with highly similar frames, we introduce STPrune to reduce redundant tokens. It consists of two key modules: Density-based Token Merging (DTM) for the segment’s visual context and Attention-based Token Selection (ATS) for salient details. In each segment, we retain $R\%$ of tokens, where $dR\%$ are obtained using DTM and $(1-d)R\%$ are obtained using ATS. $d$ controls the token allocation between the two modules.
Density-based Token Merging. For intra-segment merging, we first identify anchor tokens and to-be-merged tokens. Selecting anchor tokens from the entire segment would disrupt their spatial relationships, as anchor tokens corresponding to different components of an object might be distributed across multiple frames. To mitigate this, we restrict anchor token selection to specific frames. Specifically, we sample anchor frames at a fixed interval $p$ and select anchor tokens from these frames. The remaining tokens in the segment are treated as to-be-merged tokens. For a segment of $f$ frames, the total number of anchor frames is $\lfloor f/p\rfloor$. From each anchor frame, $dR/\lfloor f/p\rfloor\%$ of tokens are selected as anchor tokens.
Figure 4: Comparison of DTM and Cluster-based Token Merging. (a) Cluster-based methods aggregate tokens within each cluster and concatenate them, leading to a loss of positional information and disruption of the spatiotemporal structure in video data. (b) Our DTM updates anchor tokens while preserving their original positional information, maintaining structural coherence. Please refer to Table 4 for a detailed quantitative comparison.
Figure 4b illustrates DTM when the segment length is 1 and shows that it consists of three key steps. First, we follow density peaks clustering algorithms [24, 9] to compute the density score. For each token in the anchor frame $[v_{1},v_{2},\cdots,v_{N}]$, we calculate its local density $\rho_{i}$ and its distance to the closest higher-density token $\delta_{i}$, obtaining the final density score $\rho_{i}\times\delta_{i}$:
$\rho_{i}=\exp(-\frac{1}{k}\sum_{v_{j}\in\text{kNN}(v_{i})}\text{d}(v_{i},v_{j})^{2})$ (3)
$\delta_{i}=\begin{cases}\min\limits_{j:\rho_{i}>\rho_{j}}~{}\text{d}(v_{i},v_{j}),&\text{if }\exists j\text{ s.t. }\rho_{i}>\rho_{j}\\ ~{}~{}\max\limits_{j}~{}~{}\text{d}(v_{i},v_{j}),&\text{otherwise}\end{cases}$ (4)
where $\text{d}(v_{i},v_{j})$ denotes the Euclidean distance. Density peak tokens with high $\rho_{i}\times\delta_{i}$ serve as anchor tokens, indicating that they are surrounded by neighbors with lower local density while being relatively distant from other high-density tokens. This ensures anchor tokens are both representative and distinctive.
Next, Similarity-based Assignment assigns each remaining token to its nearest anchor. Finally, we apply Anchor-centric Aggregation to merge the assigned tokens into their respective anchors, preserving key visual structures through representative tokens. For an anchor $a$ and its associated tokens $[b_{1},\cdots,b_{n}]$, the updated $a^{*}$ is computed as:
$a^{*}=\beta a+\frac{1-\beta}{n}\sum_{i=1}^{n}b_{i},$ (5)
where $\beta$ controls the balance between the anchor token and its associated tokens. In Figure 4, we compare our DTM with Cluster-based Token Merging. In LLMs, RoPE [30] encodes relative positional relationships between tokens, making positional information essential for maintaining the spatiotemporal structure of video tokens. While prior methods [11] rely on Cluster-based Token Merging, they discard positional information, causing RoPE to struggle with encoding spatiotemporal structure. In contrast, our DTM maintains the positional information of merged tokens, enhancing visual context understanding. This is particularly beneficial for LLaVA-Video, which explicitly depends on positional information, where our DTM significantly outperforms other methods (see Table 4).
Overall, our DTM offers three key advantages: (1) Density-based Sampling selects density peak tokens as anchors, ensuring representative visual context. (2) We maintain the original positional information of updated anchor tokens, preserving spatiotemporal structure. (3) Anchor-centric Aggregation emphasizes representative tokens, enhancing feature representation.
Attention-based Token Selection. In addition to segment visual context obtained through DTM, we introduce ATS to capture salient visual details.
Motivated by previous studies [26, 38, 34, 41], we utilize [CLS] attention scores to identify salient visual information. However, in Video LLMs, which commonly use SigLIP [39] as their vision encoder, [CLS] attention scores cannot be obtained. This is because Video LLMs utilize the penultimate layer of SigLIP, omitting the SigLIP head where the [CLS] token is generated. To overcome this, we reintegrate a pretrained SigLIP head into the vision encoder, allowing the model to compute [CLS] attention scores. Since the SigLIP head is lightweight (15.2M parameters), this modification incurs minimal computational overhead compared to the full Video LLM.
Specifically, for a given segment, we extract the [CLS] attention score from the pretrained SigLIP head, $\mathbf{A}\in\mathbb{R}^{f\times H\times W}$, where $f$ is the number of frames, and $H$ and $W$ are the spatial dimensions of frame tokens. Since Video LLMs incorporate pooling (see Figure 2), we apply the same operation to $\mathbf{A}$, resulting in $\bar{\mathbf{A}}\in\mathbb{R}^{f\times\bar{H}\times\bar{W}}$, where $\bar{H}$ and $\bar{W}$ are the pooled spatial dimensions. Finally, for each frame, we select the top $(1-d)R\%$ of tokens with the highest attention scores to preserve critical visual details.
4 Experiments
Model Method Retention Ratio $R$ MVBench LongVideo Bench MLVU VideoMME Avg.
Overall Short Medium Long Score %
Duration 16s 1$\sim$60min 3$\sim$120min 1$\sim$60min 1$\sim$3min 3$\sim$30min 30$\sim$60min
LLaVA- OneVision [14] Vanilla 100% 56.9 56.4 65.2 58.6 70.3 56.6 48.8 59.3 100
DyCoke [31] 32% 56.3 56.6 62.1 57.1 68.1 56.7 46.7 58.0 97.8
DyCoke [31] 25% 49.5 48.1 55.8 51.0 61.1 48.6 43.2 51.1 86.2
FastV [4] 54.7 55.5 61.5 56.2 68.0 54.6 46.0 57.0 96.1
VisionZip [38] 53.7 51.2 58.5 54.1 61.6 53.4 47.2 54.4 91.7
VisionZip [38] 56.6 55.7 64.8 58.0 68.6 57.7 47.7 58.8 99.1
PruneVID [11] 55.0 55.4 64.6 57.5 69.7 54.8 48.1 58.1 98.0
FastVID 56.5 56.3 64.1 58.0 69.9 56.6 47.7 58.7 99.0
FastV [4] 20% 54.1 56.6 61.2 56.2 66.8 54.6 47.2 57.0 96.1
VisionZip [38] 53.0 50.0 57.1 53.0 60.8 51.0 47.1 53.3 90.0
VisionZip [38] 55.8 55.4 64.2 58.0 68.6 57.0 48.3 58.4 98.5
PruneVID [11] 54.6 54.7 63.7 56.2 68.1 53.9 46.7 57.3 96.6
FastVID 56.3 57.1 63.9 57.9 69.3 56.7 47.7 58.8 99.1
FastV [4] 15% 53.2 54.9 59.8 54.7 65.1 53.4 45.7 55.7 93.9
VisionZip [38] 50.3 46.9 54.4 49.5 55.8 49.3 43.3 50.3 84.8
VisionZip [38] 54.3 53.9 63.1 55.5 63.0 54.4 49.1 56.7 95.6
PruneVID [11] 54.6 53.6 63.3 56.1 67.6 53.4 47.2 56.9 96.0
FastVID 56.0 56.2 63.2 57.7 69.3 56.2 47.4 58.3 98.3
FastV [4] 10% 51.7 52.1 57.7 52.4 60.9 51.4 45.0 53.5 90.2
VisionZip [38] 44.4 43.5 51.5 46.0 50.4 45.8 41.8 46.4 78.3
VisionZip [38] 51.7 48.3 59.7 52.8 59.4 52.0 46.9 53.1 89.6
PruneVID [11] 53.2 54.4 61.7 56.0 66.6 53.0 48.4 56.3 94.9
FastVID 55.9 56.3 62.7 57.3 67.4 56.0 48.6 58.1 98.0
LLaVA- Video [43] Vanilla 100% 60.4 59.6 70.3 64.1 76.9 62.1 53.4 63.6 100
DyCoke [31] 32% 59.3 57.9 65.7 61.6 74.6 59.0 51.1 61.1 96.1
DyCoke [31] 25% 50.8 53.0 56.9 56.1 65.8 53.6 48.9 54.2 85.2
FastV [4] 58.0 58.3 63.9 61.0 71.3 60.7 51.0 60.3 94.8
VisionZip [38] 56.4 54.1 62.1 58.6 66.3 58.2 51.2 57.8 90.9
VisionZip [38] 58.3 58.3 66.6 61.9 73.3 60.1 52.2 61.5 96.7
PruneVID [11] 55.3 57.7 64.1 60.2 72.4 57.7 50.6 59.3 93.2
FastVID 59.9 57.4 68.6 63.6 74.9 62.2 53.7 62.4 98.1
FastV [4] 10% 55.8 55.4 58.9 57.9 67.6 57.7 48.6 57.0 89.6
VisionZip [38] 46.3 46.6 52.2 49.5 54.2 49.9 44.3 48.7 76.6
VisionZip [38] 56.6 53.6 61.7 58.7 67.6 58.6 50.1 57.7 90.7
PruneVID [11] 54.3 55.3 61.8 56.4 67.9 54.1 47.1 57.0 89.6
FastVID 58.5 56.5 64.9 60.7 71.7 59.1 51.2 60.2 94.7
Table 1: Comparison of state-of-the-art methods across various benchmarks. We implement the baseline methods based on their official implementations. The performance of the vanilla versions of LLaVA-OneVision and LLaVA-Video differs slightly from their reported results but remains within the margin of error. The best performance at the same retention ratio $R$ is highlighted in bold.
4.1 Experimental Settings
Benchmarks. We evaluate our method on several widely used video understanding benchmarks: MVBench [16, 25], LongVideoBench [36], MLVU [44], and VideoMME (wo sub.) [10]. Specifically, VideoMME is officially divided into short, medium, and long subsets. These benchmarks contain videos of varying durations and complex scenarios, providing a comprehensive evaluation of our method’s effectiveness and generalization.
Implementation Details. We apply our method to two representative Video LLMs: LLaVA-OneVision-7B [14] is a general Video LLM, while LLaVA-Video-7B [43] employs newline tokens based on positional information of video tokens. These models encompass different architectural characteristics, ensuring a robust evaluation of our approach.
For LLaVA-OneVision, $32$ sampled frames generate a $32\times 196$ token input to the LLM, while for LLaVA-Video, $64$ sampled frames generate a $64\times 169$ token input. Our method focuses on token compression at the LLM input stage without modifying the prefilling or decoding stages. Consequently, the retention ratio $R\%$ of input tokens directly impacts inference speed and model complexity. Unless otherwise specified, we adopt the hyperparameter setting $c=8,\tau=0.9,d=0.4,p=4,\beta=0.6$ for all experiments. We conduct all evaluations using LMMs-Eval [40].
Compared Baselines. We compare FastVID with strong baselines from both image and video compression methods.
For image compression, we adopt both the widely used classic method FastV [4] and the current state-of-the-art approach VisionZip [38]. FastV prunes tokens during the prefilling stage based on attention scores from query tokens. In contrast, VisionZip prunes tokens at the input stage. However, VisionZip prunes tokens in the last layer of the vision encoder, conflicting with pooling in Video LLMs and degrading performance. To address this, we implement VisionZip, which applies pruning after pooling.
For video compression, we compare two of the latest methods, DyCoke [31] and PruneVID [11]. Both employ multi-stage pruning across the input, prefilling, and decoding stages. Although they initially prune redundant video tokens at the input stage, their prefilling and decoding processes disregard video-specific spatiotemporal redundancy. Instead, they treat video tokens as textual tokens and prune them solely based on attention scores from query tokens, similar to FastV. Furthermore, their input-stage video pruning remains inefficient, retaining $47.5\%$ of tokens in DyCoke and $50\%$ in PruneVID. This indicates that their overall efficiency is primarily influenced by prefilling and decoding stages rather than effective video redundancy removal. Since our study focuses on video spatiotemporal redundancy, we ensure a fair comparison by evaluating only their video pruning strategies.
Segmentation MVBench LongVideo- Bench MLVU VideoMME Avg.
Fixed-interval 55.1 53.6 61.7 55.4 95.3
Cluster-based 53.2 53.0 61.7 54.8 93.9
Our DySeg 55.9 56.3 62.7 57.3 98.0
Table 2: Ablation study on different video segmentation methods. As shown in Figure 3, Fixed-interval Segmentation partitions the video at equal intervals. Cluster-based Segmentation groups frames based on frame similarity. Our proposed DySeg adaptively segments the video based on transition similarity.
4.2 Comparisons with State-of-the-Art Methods
For a comprehensive evaluation, Table 1 compares our FastVID with state-of-the-art methods on benchmarks with diverse video durations. In DyCoke [31], video frames are segmented into groups of 4, retaining all tokens from the first frame while pruning the rest. Consequently, its lowest retention ratio $R$ is $25\%$. For other baselines, we conduct comparisons across different $R$ values.
Results on LLaVA-OneVision. We evaluate our FastVID against other methods on LLaVA-OneVision at retention ratios of $25\%,20\%,15\%,10\%$, comparing their token compression effectiveness. Image compression methods perform well at $R=25\%,$ but their performance declines sharply as $R$ decreases. For instance, from $R=25\%$ to $R=10\%$, FastV and VisionZip exhibit performance drops of $5.9\%$ and $9.5\%$, respectively. This indicates the limitations of spatial compression alone, which struggles to preserve essential temporal information under extreme pruning. When $R=10\%$, video compression methods, particularly PruneVID and our FastVID, significantly outperform image-based methods, demonstrating the importance of modeling temporal redundancy for efficient video understanding. While our FastVID achieves performance comparable to VisionZip at $R=25\%$, its advantage becomes increasingly evident as $R$ decreases, with an average performance drop of just $\textbf{1.0}\%$. Notably, even after pruning $\textbf{90}\%$ of the tokens, our FastVID preserves $\textbf{98.0}\%$ of the vanilla model’s performance. FastVID significantly outperforms PruneVID by $\textbf{3.1}\%$, FastV by $\textbf{7.8}\%$, and VisionZip by $\textbf{8.4}\%$, demonstrating an optimal trade-off between efficiency and accuracy.
$d$ in STPrune MVBench LongVideo- Bench MLVU VideoMME Avg.
0.0/ATS 55.3 55.3 62.4 55.0 96.2
0.2 55.0 53.6 62.1 56.1 95.7
0.4 55.9 56.3 62.7 57.3 98.0
0.6 55.6 54.5 62.0 56.2 96.3
0.8 56.0 54.9 62.6 55.6 96.7
1.0/DTM 54.1 50.1 61.5 55.9 93.5
Table 3: Ablation Study on $d$ in STPrune. STPrune consists of DTM and ATS. A fraction $d$ of the pruned tokens is from DTM, while the remaining $1-d$ is from ATS. When $d=0.0$, all pruned tokens are from ATS, while at $d=1.0$, they are from DTM.
Results on LLaVA-Video. We conduct experiments on LLaVA-Video with $R=25\%$ and $R=10\%$. When $R=10\%$, the performance of both video compression methods in LLaVA-Video is lower than in LLaVA-OneVision. This discrepancy is likely due to LLaVA-Video’s higher pooling compression rate, where each frame contains $169$ tokens instead of $196$ tokens in LLaVA-OneVision. Further pruning under this higher pooling compression rate leads to a performance drop. When $R=25\%$, image compression methods achieve state-of-the-art performance on LongVideoBench. However, FastVID outperforms them significantly on other datasets, achieving superior $\textbf{98.1}\%$ performance. When $R=10\%$, FastVID consistently surpasses all other methods across all benchmarks. On LLaVA-Video, our FastVID maintains a clear advantage, outperforming other methods by $\textbf{1.4}\%$ at $R=25\%$ and $\textbf{4.0}\%$ at $R=10\%$.
4.3 Ablation Study
In this section, we conduct ablation studies on the proposed modules and hyperparameters. By default, we perform experiments on LLaVA-OneVision, setting the retention ratio $R$ to $10\%$. Under this configuration, the $32\times 196$ video tokens are compressed to an average of $19$ tokens per frame.
Ablation study on DySeg. To compare different video segmentation methods, we present qualitative and quantitative results in Figure 3 and Table 2, respectively. In Table 2, Fixed-interval Segmentation with an interval of $4$ generates $32/4=8$ segments, while Cluster-based Segmentation generates $8$ clusters, ensuring consistency with the minimal segment number $c$ in our DySeg. Fixed-interval preserves temporal order between segments, whereas Cluster-based maintains intra-segment similarity. The results show that Fixed-interval outperforms Cluster-based by $1.4\%$, highlighting the importance of temporal structure preservation in Video LLM pruning. Additionally, achieving high intra-segment similarity is also essential for efficient pruning. Our proposed DySeg successfully integrates both advantages while enabling dynamic pruning strategies, achieving a superior average performance of $\textbf{98.0}\%$.
Token Merging MVBench LongVideo- Bench MLVU VideoMME Avg.
LLaVA-OneVision
Uniform 55.0 54.7 62.0 56.4 96.2
Cluster-based 55.6 55.2 62.4 57.3 97.2
Our DTM 55.9 56.3 62.7 57.3 98.0
LLaVA-Video
Uniform 55.4 54.6 62.9 59.8 91.5
Cluster-based 55.3 55.1 62.9 60.1 91.8
Our DTM 58.5 56.5 64.9 60.7 94.7
Table 4: Ablation study on different token merging strategies. Uniform, adopted in VisionZip (see Figure 1c), merges tokens through uniform sampling. Cluster-based methods (see Figure 4a) conduct clustering algorithms. Our proposed DTM employs a density-based method (see Figure 4b).
Ablation study on STPrune. Our STPrune is proposed to prune tokens within each segment and consists of two key components: DTM and ATS. DTM employs density-based sampling to merge redundant tokens, preserving segment visual context. ATS leverages [CLS] attention scores to highlight important visual details. Table 3 presents an ablation study on $d$, which controls the token pruning distribution between DTM and ATS. The results show that ATS alone (when $d=0.0$) outperforms DTM alone (when $d=1.0$). However, the best performance is achieved at $d=0.4$, demonstrating that a balanced integration of both modules is crucial for preserving essential video information. This integration boosts performance by $\textbf{1.8}\%$ and $\textbf{4.5}\%$ points over ATS and DTM alone, respectively.
Ablation study on DTM. In Table 4, we compare different token merging strategies for segment visual context. Our DTM consistently outperforms other approaches in both Video LLMs. Uniform struggles to distinguish objects due to indiscriminate anchor token selection (see Figure 1c). Cluster-based discards essential positional information (see Figure 4a), resulting in inferior performance, particularly in LLaVA-Video. By leveraging Density-based Sampling and Anchor-centric Aggregation, our DTM achieves superior results, especially in LLaVA-Video, where DTM outperforms other methods by $\textbf{2.9}\%$.
Ablation study on $c$ and $\tau$ in DySeg. We evaluate the effect of $c$ and $\tau$ in Eq. (2) of DySeg. Figure 5(a) shows results with varing $c$. In simple video scenarios with high transition similarity, $c$ regulates segmentation. When $c=1$, segmentation relies solely on $\tau$. $c=32$ divides the video into single-frame segments. Performance remains stable for $1\leq c\leq 16$, but the decline at $c=32$ suggests that excessive segmentation disregards temporal relationships. The optimal performance at $c=8$ indicates the benefit of a minimum cluster number. Figure 5(b) shows results with varing $\tau$. When $\tau$ is small, most transitions are selected by $c$. As $\tau$ increases, more transitions fall below the threshold. The best performance is achieved at $\tau=0.9$, effectively grouping redundant frames while separating non-redundant ones.
Ablation study on $p$ and $\beta$ in DTM. In Figure 6(a), we evaluate the effect of $p$ on anchor frame selection. Anchor tokens are sampled from anchor frames. When $p=1$, anchor tokens are evenly distributed across all frames within a segment. As $p$ increases, tokens concentrate on fewer frames. Given the high similarity between frames in a segment, a sufficient number of anchor tokens per anchor frame is essential to capture spatial context. However, if $p$ is too large, all anchor tokens come from the first frame, limiting temporal information. We find that $p=4$ yields the best performance by effectively capturing spatiotemporal context. Figure 6(b) shows the effect of $\beta$ in Eq. (5). When $\beta=0.0$, matching tokens are averaged. $\beta=1.0$ discards all non-anchor tokens, removing segment visual context and leading to a performance drop. Notably, Anchor-centric Assignment (when $\beta=0.6$) yields optimal results, highlighting the importance of representative tokens and visual context.
Figure 5: Ablation study on $c$ and $\tau$ in DySeg. $c$ denotes the minimum number of segments in a video, whereas $\tau$ denotes the transition similarity threshold. Both parameters jointly control the segmentation granularity.
Figure 6: Ablation study on $p$ and $\beta$ in DTM. $p$ denotes the interval for anchor frame selection, whereas $\beta$ controls the merging weight for anchor tokens and their associated tokens in Eq. (5).
5 Conclusion
In this paper, we introduced FastVID, a novel inference-time pruning framework designed to accelerate Video LLMs by effectively reducing spatiotemporal redundancy. Through a comprehensive analysis of video tokens from both temporal context and visual context, FastVID dynamically partitions videos into temporally ordered segments and employs density-based token pruning within each segment. This approach significantly decreases computational overhead while maintaining crucial temporal and visual integrity. Extensive experiments across different Video LLMs and diverse video benchmarks confirm the effectiveness and strong generalization capability of our method.