### v2

# FastVID: Dynamic Density Pruning for Fast Video Large Language Models

Abstract
[…]
1 Introduction
Video Large Language Models (Video LLMs) [6, 19, 16, 47, 37] have shown strong performance in video understanding but incur substantial inference costs. While several methods [31, 24, 40, 15, 30, 49] explore training-time compression to mitigate this issue, they often require expensive retraining. In this work, we focus on an inference-time acceleration strategy that enhances efficiency without requiring additional training.
Figure 1: Effective token compression in Video LLMs relies on preserving both temporal and visual context. (a) shows the impact of temporal disruptions, emphasizing the importance of temporal structure preservation. (b) show the effects of different token selection strategies for visual merging, where patches that share the same inner and border color are merged. Density-based Sampling retains both distinctive and representative context while effectively reducing redundancy.
This computational burden is primarily caused by the high volume of video tokens, making effective token compression essential. While prior image compression methods [5, 28, 46, 42, 4] reduce redundancy within a single image, they fail to exploit temporal dependencies across frames. As a result, the nature of video redundancy remains underexplored. In this work, we systematically analyze video redundancy from two key perspectives: temporal context and visual context.
Temporal context is fundamental to video understanding, as the order and continuity of frames directly influence semantic interpretation. As depicted in Figure 1(a), disrupting frame order (shuffled) or omitting frames (incomplete) leads to incorrect comprehension, highlighting the necessity of preserving temporal structure. To achieve this, we segment the video into temporally ordered segments, grouping highly similar consecutive frames. Pruning is applied within each high-redundancy segment but not across segments, preserving the essential temporal structure for accurate video understanding.
To maintain visual context, redundant tokens within each segment must be effectively merged. A state-of-the-art method VisionZip [42] applies uniform token sampling followed by merging redundant tokens [4] as shown in Figure 1(b1). However, its token selection strategy is content-agnostic, potentially leading to the loss of important details, such as the guitar body being incorrectly merged with the background. To address this, we propose a density-based token sampling in Figure 1(b2). Specifically, high-density tokens, surrounded by numerous similar tokens, serve as candidates for selection. We select density peak tokens as anchors, which have a higher density than their neighbors and by a relatively large distance from tokens with higher densities [26]. This strategy ensures that the selected tokens are both representative and distinctive, effectively preserving segment visual context while reducing redundancy.
Building upon these insights, we propose Dynamic Density Pruning for Fast VIDeo LLMs, minimizing spatiotemporal redundancy while preserving essential semantics. We begin with Dynamic Temporal Segmentation, which adaptively partitions videos into segments. Within each segment, we introduce Density Spatiotemporal Pruning to retain both global visual context and salient details. FastVID significantly accelerates inference while preserving both temporal and visual integrity.
[…]
The main contributions are summarized as follows: (1) We analyze Video LLM compression from both temporal and visual context perspectives, emphasizing the importance of maintaining temporal and visual integrity. (2) We propose FastVID, a novel pruning framework that employs Dynamic Temporal Segmentation to partition videos into temporally ordered segments and Density Spatiotemporal Pruning to retain global segment information and key details. (3) Our FastVID achieves state-of-the-art performance across diverse video benchmarks and maintains robust accuracy even under extreme compression.
2 Related Work
Video LLMs. With the rapid advancement of LMMs [1, 7, 34] and MLLMs [35, 21, 22, 17, 3, 23], there has been growing interest in Video LLMs. These models can be categorized based on how they process video tokens: general Video LLMs and Video LLMs with training-time compression.
General Video LLMs [20, 6, 19, 16, 47, 37] directly process raw video tokens or apply pooling. Video-LLaVA [19] leverages shared projection layers to obtain unified visual representations. LLaVA-OneVision [16] demonstrates strong video understanding through task transfer from images. LLaVA-Video [47] creates a high-quality synthetic dataset for video instruction-following. To better capture the spatiotemporal structure of video, some models [47, 37] introduce additional designs for video positional information. LLaVA-Video [^1] https://github.com/LLaVA-VL/LLaVA-NeXT introduces newline tokens to distinguish spatial and temporal positions effectively.
Video LLMs with training-time compression [31, 24, 40, 15, 30, 49] aim to significantly reduce the number of video tokens, enabling longer video sequences. PLLaVA [40] introduces an adaptive average structure pooling to extend image LLMs. Video-ChatGPT [24] extracts both spatial and temporal features through temporal and spatial pooling respectively. Chat-UniVi [15] progressively clusters visual tokens and provides multi-scale features. LongVU [30] employs cross-modal query and inter-frame dependencies to adaptively reduce video redundancy. Apollo [49] explores scaling consistency and uses the Perceiver Resampler [14].
However, general Video LLMs remain the dominant paradigm, with LLaVA-OneVision being widely adopted due to its adaptability and superior performance. Therefore, we focus on inference-time acceleration for general Video LLMs. To evaluate our method, we select LLaVA-OneVision and LLaVA-Video as representative models.
Token Compression. Token compression has emerged as an effective approach to reduce computational complexity in transformer architectures, such as ViT [9] and CLIP [25]. ToMe [4] progressively merges fixed spatial tokens, while TempMe [29] extends this concept by merging neighboring clips to minimize temporal redundancy.
Recent studies [5, 28, 46, 42, 4, 39] primarily focus on spatial token reduction to accelerate Image LLMs. FastV [5] selects text-relevant tokens at shallow layers of the LLM. LLaVA-PruMerge [28] uses attention scores from the [CLS] token to prune spatial redundancy. SparseVLM [46] proposes a token recycling strategy to aggregate and reconstruct tokens to be pruned. VisionZip [42] reduces visual redundancy in the vision encoders. However, these methods overlook the temporal relationships across frames.
Due to the high volume of video tokens in Video LLMs, recent video compression methods [12, 33, 13] have gained increasing attention. FrameFusion [12] applies both merging and pruning across successive shallow LLM layers, but repeated pruning operations adversely affect overall efficiency. DyCoke [33] merges tokens across frames and applies dynamic KV cache reduction. However, its pruning during the prefilling stage struggles to achieve substantial token reduction while maintaining accuracy. PruneVID [13] clusters video tokens and selects those most relevant to query tokens, but this dependency on clustering introduces significant latency during compression. Despite these advances, efficient and accurate pruning under large token reduction remains unsolved.
In this paper, we focus on preserving temporal and visual context for spatiotemporal token pruning at the prefilling stage. By reducing video tokens before LLM processing, our FastVID significantly enhances computational efficiency while facilitating easy deployment, including compatibility with FlashAttention [8], KV cache, and a plug-and-play design for seamless integration into existing Video LLMs. Our method achieves robust performance even under extreme compression rates, offering a practical solution for fast Video LLMs.
3 Methodology
3.1 FastVID
Video LLMs face significant inference challenges due to the large number of frames in a video and the numerous vision tokens within each frame. Typically, the total token count often exceeds the LLM’s context length, making direct processing computationally infeasible. Although uniform frame sampling and token pooling methods are commonly employed to reduce complexity, they still leave a substantial number of redundant video tokens (e.g., $32\times 196=6272$ video tokens in LLaVA-OneVision [16]).
To address these limitations, we propose FastVID, a novel method that accelerates inference by reducing redundant tokens while preserving critical temporal and visual context. First, we introduce Dynamic Temporal Segmentation (DySeg) in Section 3.2, which partitions a video into temporally ordered segments with high redundancy. Within each segment, we introduce Density Spatiotemporal Pruning (STPrune) in Section 3.3, which retains essential visual context and salient details.
Figure 2: The overview of our FastVID.
Figure 2 presents an overview of FastVID. Given an input video, $F$ frames are uniformly sampled (e.g., 32 in LLaVA-OneVision, 64 in LLaVA-Video). Each frame is individually processed by the vision encoder. The extracted tokens are projected and pooled into a video token sequence. With FastVID, we effectively eliminate redundant tokens while preserving critical information. Specifically, DySeg dynamically partitions video tokens into temporally ordered, high-redundancy segments, while STPrune performs density-based pruning within each segment, enhancing efficiency without loss of key information. The final retained video tokens, combined with query tokens, is then fed into the LLM to generate the response.
3.2 Dynamic Temporal Segmentation
As shown in Figure 3(b) and 3(c), there are two common static segmentation methods for sampled frame sequences. In Figure 3(b), Fixed-interval Segmentation partitions the sequence into segments of a fixed length, preserving temporal order but potentially grouping visually dissimilar frames. Figure 3(c) shows Cluster-based Segmentation, where frames are grouped into three clusters based on frame similarity. However, it suffers from a predefined cluster number, leading to ineffective segmentation when video complexity varies. As a result, the first segment contains similar objects but different scenes. Furthermore, clustering may disrupt the temporal order by ignoring critical temporal information, such as omitting key frames (e.g., the 9th frame) and incorrectly grouping frames from different time periods into the first segment.
To address the limitations of static methods, we propose Dynamic Temporal Segmentation, a simple yet effective method that adaptively refines segment boundaries according to video complexity. DySeg achieves both temporal structure and high intra-segment similarity, generating fewer partitions for simple scenes and finer ones for more complex scenes.
To enable effective pruning, DySeg promotes high spatiotemporal redundancy within each segment by minimizing similarity between adjacent segments. Thus, we segment the video based on transition similarity between adjacent frames. Specifically, we utilize global frame features $\mathbf{f}$ to compute the cosine similarity:
$t_{i}=\cos(\mathbf{f}_{i},\mathbf{f}_{i+1}),\quad i=1,\cdots,F-1,$
$\mathbf{T}=\{t_{1},t_{2},\cdots,t_{F-1}\},$ (1)
where $t_{i}$ denotes the transition similarity between the $i$-th and $(i+1)$-th frame. $\mathbf{T}$ denotes the $F-1$ transition similarities for $F$ sampled frames. To achieve fine-grained segmentation, we select transitions that satisfy the following conditions:
$\mathbf{S}_{1}=\operatorname{arg\,min}_{c-1}\mathbf{T},\quad\mathbf{S}_{2}=\{i\mid t_{i}<\tau,\;t_{i}\in\mathbf{T}\},$
$\mathbf{S}=\mathbf{S}_{1}\cup\mathbf{S}_{2},$ (2)
where $c$ denotes the minimum number of segments and $\tau$ denotes the threshold for transition similarity. $\mathbf{S}_{1}$ denotes the $c-1$ least similar frame transitions, while $\mathbf{S}_{2}$ denotes transitions where similarity falls below a fixed threshold $\tau$. Each transition in the union $\mathbf{S}$ marks a boundary between segments. In simple videos with minimal shot transitions, $\mathbf{S}_{2}$ is often empty. In such cases, $\mathbf{S}_{1}$ enables finer segmentation by distinguishing subtle temporal changes. For complex videos, $\mathbf{S}_{1}$ is typically a subset of $\mathbf{S}_{2}$, where $\mathbf{S}_{2}$ ensures that adjacent frames with similarity below $\tau$ are assigned to different segments. In Figure 3(d), our DySeg effectively segments videos by dynamically adjusting granularity in a simple yet effective manner, outperforming Fixed-interval and Clustering-based methods.
3.3 Density Spatiotemporal Pruning
After obtaining segments with highly similar frames, we introduce STPrune to reduce redundant tokens. It consists of two key modules: Density-based Token Merging (DTM) for segment visual context and Attention-based Token Selection (ATS) for salient details. For a segment of $P$ frames, we retain $rPN$ tokens in total, where $N$ is the number of tokens per frame, and $r$ denotes the retention ratio. These tokens are generated by DTM and ATS. Specifically, $drFN$ tokens are generated by DTM and $(1-d)rFN$ tokens are generated by ATS, where $d$ controls the token allocation between the two modules.
Density-based Token Merging. For segment visual context, we first identify anchor tokens and to-be-merged tokens. Selecting anchor tokens from the entire segment would disrupt their spatial relationships, as anchor tokens corresponding to different components of the same object might be distributed across multiple frames. To mitigate this, we restrict anchor token selection to specific frames. Specifically, we sample anchor frames at a fixed interval $p$ and select anchor tokens from these frames. The remaining tokens in the segment are treated as to-be-merged tokens. Specifically, the number of anchor frames is $\lceil P/p\rceil$. From each anchor frame, $drPN/\lceil P/p\rceil$ tokens are selected as anchor tokens. Please see Figure 5 for a visualization of DTM applied to video segments.
Figure 3: Visualization of different segmentation methods on 12 sampled frames from a video example. (b) Fixed-interval Segmentation struggles to maintain high intra-segment similarity, leading to visually diverse frames within the same segment. (c) Cluster-based Segmentation disrupts temporal order, grouping frames from different time periods to the same segment. (d) Our DySeg adaptively partitions the video, preserving both temporal structure and high intra-segment similarity. Please refer to Table 6 for a detailed quantitative comparison.
To further clarify DTM, Figure 4(b) illustrates DTM when the segment length is 1 and shows that it consists of three key steps. First, we follow density peaks clustering algorithms [26, 10] to compute the density score. For each token in the anchor frame $[v_{1},v_{2},\cdots,v_{N}]$, we calculate its local density $\rho_{i}$ and its distance to the closest higher-density token $\delta_{i}$, obtaining the final density score $\rho_{i}\times\delta_{i}$:
$\rho_{i}=\exp(-\frac{1}{k}\sum_{v_{j}\in\text{kNN}(v_{i})}\text{d}(v_{i},v_{j})^{2}),$ (3)
$\delta_{i}=\begin{cases}\min\limits_{j:\rho_{j}>\rho_{i}}~{}\text{d}(v_{i},v_{j}),&\text{if }\exists j\text{ s.t. }\rho_{j}>\rho_{i}\\ ~{}~{}\max\limits_{j}~{}~{}\text{d}(v_{i},v_{j}),&\text{otherwise}\end{cases}$ (4)
where $\text{d}(v_{i},v_{j})$ denotes the Euclidean distance. Density peak tokens with high $\rho_{i}\times\delta_{i}$ serve as anchor tokens, indicating that they are surrounded by neighbors with lower local density while remaining relatively distant from other high-density tokens. This selection ensures that anchor tokens are both representative and distinctive. Next, for each anchor frame, Similarity-based Assignment assigns each token in the segment to the nearest anchor token using cosine similarity. Finally, we apply Anchor-centric Aggregation to merge the assigned tokens into their respective anchors, preserving key visual structures through representative tokens. For an anchor $a$ and its associated tokens $[b_{1},\cdots,b_{n}]$, the updated $a^{*}$ is computed as:
$a^{*}=\alpha a+\frac{1-\alpha}{n}\sum_{i=1}^{n}b_{i},$ (5)
where $\alpha$ controls the balance between the anchor token and its associated tokens. In Figure 4, we compare our DTM with Cluster-based Token Merging. In LLMs, RoPE [32] encodes relative positional relationships between tokens, making positional information essential for maintaining the spatiotemporal structure of video tokens. While prior methods [13] rely on Cluster-based Token Merging, they discard positional information, causing RoPE to struggle with encoding spatiotemporal structure. In contrast, our DTM maintains the positional information of merged tokens, enhancing visual context understanding.
Overall, our DTM offers three key advantages: (1) Density-based Sampling selects density peak tokens as anchors, ensuring representative visual context. (2) We maintain the original positional information of updated anchor tokens, preserving spatiotemporal structure. (3) Anchor-centric Aggregation emphasizes representative tokens, enhancing feature representation.
Figure 4: Comparison of DTM and Cluster-based Token Merging. (a) Cluster-based methods aggregate tokens within each cluster and concatenate them, leading to a loss of positional information and disruption of the spatiotemporal structure in video data. (b) Our DTM updates anchor tokens while preserving their original positional information, maintaining structural coherence. Please refer to Table 5 for a detailed quantitative comparison.
Attention-based Token Selection. In addition to segment visual context obtained through DTM, we introduce ATS to capture salient visual details.
Motivated by previous studies [28, 42, 36, 45], we utilize [CLS] attention scores to identify salient visual information. However, in Video LLMs, which commonly use SigLIP [43] as their vision encoder, [CLS] attention scores cannot be obtained. This is because Video LLMs utilize the penultimate layer of SigLIP, omitting the SigLIP head where the [CLS] token is generated. To overcome this, we reintegrate a pretrained SigLIP head into the vision encoder, allowing the model to compute [CLS] attention scores. Since the SigLIP head is lightweight (15.2M parameters), this modification incurs minimal computational overhead compared to the full Video LLM.
Specifically, we extract the [CLS] attention score from the pretrained SigLIP head, $\mathbf{A}\in\mathbb{R}^{H\times W}$, where $H$ and $W$ are the spatial dimensions of frame tokens. Since Video LLMs incorporate pooling (see Figure 2), we apply the same operation to $\mathbf{A}$, resulting in $\bar{\mathbf{A}}\in\mathbb{R}^{\bar{H}\times\bar{W}}$, where $\bar{H}$ and $\bar{W}$ are the pooled spatial dimensions. Finally, for each frame, we select the top $(1-d)rN$ tokens with the highest attention scores to preserve critical visual details.
4 Experiments
4.1 Experimental Settings
Benchmarks. We evaluate our method on several widely used video understanding benchmarks: MVBench [18, 27], LongVideoBench [38], MLVU [48], and VideoMME (wo sub.) [11]. Specifically, VideoMME is officially divided into short, medium, and long subsets. These benchmarks contain videos of varying durations and complex scenarios, providing a comprehensive evaluation of our method’s effectiveness and generalization.
Table 1: Comparison of state-of-the-art methods on LLaVA-OneVision [16]. The A%/Bretention ratio indicates that Aof the LLM input tokens are retained, and subsequently compressed to Bduring the LLM forward pass. The best performance among those with similar retention ratios $R$ is highlighted in bold. TFLOPs related to video tokens are reported (see the appendix for details).
Method Retention Ratio $R$ TFLOPs MVBench LongVideo Bench MLVU VideoMME Avg. Acc.
Overall Short Medium Long Score %
Duration 16s 1$\sim$60min 3$\sim$120min 1$\sim$60min 1$\sim$3min 3$\sim$30min 30$\sim$60min
Vanilla 100% 48.82 56.9 56.4 65.2 58.6 70.3 56.6 48.8 59.3 100
DyCoke [33]${}_{\text{CVPR'25}}$ 32.5% 14.13 56.3 56.6 62.1 57.1 68.1 56.7 46.7 58.0 97.8
FastV [5]${}_{\text{ECCV'24}}$ 100%/25% 13.45 54.7 55.5 61.5 56.2 68.0 54.6 46.0 57.0 96.1
VisionZip [42]${}_{\text{CVPR'25}}$ 25% 10.73 53.7 51.2 58.5 54.1 61.6 53.4 47.2 54.4 91.7
VisionZip [42]${}_{\text{CVPR'25}}$ 25% 10.73 56.6 55.7 64.8 58.0 68.6 57.7 47.7 58.8 99.1
DyCoke [33]${}_{\text{CVPR'25}}$ 25% 10.73 49.5 48.1 55.8 51.0 61.1 48.6 43.2 51.1 86.2
FastVID 25% 10.73 56.5 56.3 64.1 58.0 69.9 56.6 47.7 58.7 99.0
FastV [5]${}_{\text{ECCV'24}}$ 100%/20% 11.38 54.1 56.6 61.2 56.2 66.8 54.6 47.2 57.0 96.1
VisionZip [42]${}_{\text{CVPR'25}}$ 19.9% 8.46 53.0 50.0 57.1 53.0 60.8 51.0 47.1 53.3 90.0
VisionZip [42]${}_{\text{CVPR'25}}$ 19.9% 8.46 55.8 55.4 64.2 58.0 68.6 57.0 48.3 58.4 98.5
FastVID 19.9% 8.46 56.3 57.1 63.9 57.9 69.3 56.7 47.7 58.8 99.1
FastV [5]${}_{\text{ECCV'24}}$ 100%/15% 9.35 53.2 54.9 59.8 54.7 65.1 53.4 45.7 55.7 93.9
VisionZip [42]${}_{\text{CVPR'25}}$ 14.8% 6.23 50.3 46.9 54.4 49.5 55.8 49.3 43.3 50.3 84.8
VisionZip [42]${}_{\text{CVPR'25}}$ 14.8% 6.23 54.3 53.9 63.1 55.5 63.0 54.4 49.1 56.7 95.6
FastVID 14.8% 6.23 56.0 56.2 63.2 57.7 69.3 56.2 47.4 58.3 98.3
FastV [5]${}_{\text{ECCV'24}}$ 100%/10% 7.36 51.7 52.1 57.7 52.4 60.9 51.4 45.0 53.5 90.2
VisionZip [42]${}_{\text{CVPR'25}}$ 9.7% 4.04 44.4 43.5 51.5 46.0 50.4 45.8 41.8 46.4 78.3
VisionZip [42]${}_{\text{CVPR'25}}$ 9.7% 4.04 51.7 48.3 59.7 52.8 59.4 52.0 46.9 53.1 89.6
PruneVID [13]${}_{\text{ACL'25}}$ 10.1% 4.23 54.2 53.8 62.3 55.9 66.4 52.9 48.3 56.6 95.4
FastVID 9.7% 4.04 55.9 56.3 62.7 57.3 67.4 56.0 48.6 58.1 98.0
PruneVID [13]${}_{\text{ACL'25}}$ 10.1%/3.9% 2.58 54.1 51.8 62.3 55.5 67.1 51.8 47.7 55.9 94.3
FastVID+FastV [5] 9.7%/3.6% 2.47 55.0 53.6 61.9 56.3 66.1 54.8 48.1 56.7 95.6
Table 2: Comparison of state-of-the-art methods on LLaVA-Video [47]. TFLOPs calculations include both video and newline tokens.
Method Retention Ratio $R$ Newline Tokens $M$ TFLOPs MVBench LongVideo Bench MLVU VideoMME Avg. Acc.
Overall Short Long Score %
Vanilla 100% 832 103.2 60.4 59.6 70.3 64.1 76.9 53.4 63.6 100
DyCoke [33]${}_{\text{CVPR'25}}$ 32.1% 256 27.1 59.3 57.9 65.7 61.6 74.6 51.1 61.1 96.1
FastV [5]${}_{\text{ECCV'24}}$ 100%/25% 832/568.7 29.2 58.0 58.3 63.9 61.0 71.3 51.0 60.3 94.8
VisionZip [42]${}_{\text{CVPR'25}}$ 24.9% 64 19.5 56.4 54.1 62.1 58.6 66.3 51.2 57.8 90.9
VisionZip [42]${}_{\text{CVPR'25}}$ 24.9% 64 19.5 58.3 58.3 66.6 61.9 73.3 52.2 61.5 96.7
DyCoke [33]${}_{\text{CVPR'25}}$ 25% 208 20.7 50.8 53.0 56.9 56.1 65.8 48.9 54.2 85.2
FastVID 24.9% 64 19.5 59.3 58.3 67.7 62.6 74.9 52.0 62.0 97.5
FastVID 24.9% 715.5 24.5 59.9 57.4 68.6 63.6 74.9 53.7 62.4 98.1
FastV [5]${}_{\text{ECCV'24}}$ 100%/10% 832/311.8 16.2 55.8 55.4 58.9 57.9 67.6 48.6 57.0 89.6
VisionZip [42]${}_{\text{CVPR'25}}$ 9.5% 64 7.3 46.3 46.6 52.2 49.5 54.2 44.3 48.7 76.6
VisionZip [42]${}_{\text{CVPR'25}}$ 9.5% 64 7.3 56.6 53.6 61.7 58.7 67.6 50.1 57.7 90.7
FastVID 9.5% 64 7.3 58.3 56.2 63.9 59.6 70.9 50.7 59.5 93.6
FastVID 9.5% 508.2 10.5 58.5 56.5 64.9 60.7 71.7 51.2 60.2 94.7
Table 3: Efficiency Comparison on LLaVA-OneVision [16]. The prefilling time, defined as the latency to the first generated token, is measured on VideoMME using an A100 GPU.
Method # Token TFLOPs Prefilling Time (ms) Avg. Acc.
Segmentation Compression LLM Forward Total
Vanilla 6272 (100%) 48.82 (100%) 476.3 476.3 (1.0$\times$) 59.3 (100%)
PruneVID [13] 635.6 (10.1%) 4.23 (8.7%) 5.2 32.0 64.3 101.5 (4.7$\times$) 56.6 (95.4%)
FastVID 608 (9.7%) 4.04 (8.3%) 0.5 5.6 61.1 67.2 (7.1$\times$) 58.1 (98.0%)
Implementation Details. We apply our method to two representative Video LLMs: LLaVA-OneVision-7B [16] is a general Video LLM, while LLaVA-Video-7B [47] employs newline tokens based on positional information of video tokens. These models encompass different architectural characteristics, ensuring a robust evaluation of our approach. Unless otherwise specified, we adopt the hyperparameter setting $c=8,\tau=0.9,d=0.4,p=4,\alpha=0.6$ for all experiments. For LLaVA-OneVision, $32$ sampled frames generate a $32\times 196$ token input to the LLM. We test with $r\in\{25\%,20\%,15\%,10\%\}$. For LLaVA-Video, $64$ sampled frames generate a $64\times 169$ token input. We test with $r\in\{25\%,10\%\}$. We conduct all evaluations using LMMs-Eval [44] on A100 GPUs.
Compared Baselines. (1) For image compression, we adopt both the widely used classic method FastV [5] and the current state-of-the-art approach VisionZip [42]. VisionZip prunes tokens in the last layer of the vision encoder, conflicting with pooling in Video LLMs and degrading performance. To address this, we implement VisionZip, which applies pruning after pooling. (2) For video compression, we compare two recent methods, DyCoke [33] and PruneVID [13]. Our FastVID focuses on pruning during the prefilling stage. To ensure fairness, we reimplement these baselines without pruning in the decoding stage. PruneVID applies two-stage pruning on both input tokens and the LLM’s 10th layer. PruneVID refers to the variant that prunes only input tokens. More implementation details of these baselines are provided in the appendix.
4.2 Comparisons with State-of-the-Art Methods
For a comprehensive evaluation, we compare our FastVID with state-of-the-art methods on benchmarks with diverse video durations. In DyCoke, video frames are evenly divided into 4-frame segments, retaining all tokens from the first frame while pruning the rest. Consequently, its lowest retention ratio $R$ is $25\%$. For other baselines, we conduct comparisons across different $R$ values.
Results on LLaVA-OneVision. In Table 1, we evaluate our FastVID against other methods on LLaVA-OneVision. While FastV and VisionZip perform well at $R=25\%$, but their performance declines sharply as $R$ decreases. This indicates the limitations of spatial compression alone, which struggles to preserve essential temporal information under extreme pruning. Notably, even after pruning 90.3% of the tokens, our FastVID preserves 98.0% of the vanilla model’s performance. To compare with PrunVID, we additionally apply FastV at the 10th layer of the LLM. FastVID+FastV achieves 95.6% of the original accuracy at 2.47 (5.1%) TFLOPs.
Table 4: Ablation Study on $d$ in STPrune. STPrune consists of DTM and ATS. A fraction $d$ of the retained tokens is from DTM, while the remaining $1-d$ is from ATS.
$d$ in STPrune MVBench LongVideo- Bench MLVU VideoMME Avg. Acc.
0.0/ATS 55.3 55.3 62.4 55.0 96.2
0.2 55.0 53.6 62.1 56.1 95.7
0.4 55.9 56.3 62.7 57.3 98.0
0.6 55.6 54.5 62.0 56.2 96.3
0.8 56.0 54.9 62.6 55.6 96.7
1.0/DTM 54.1 50.1 61.5 55.9 93.5
Table 5: Ablation study on different token merging strategies. Figure 1(b1) presents Uniform. Figure 4(a) presents Cluster-based.
Token Merging MVBench LongVideo- Bench MLVU VideoMME Avg. Acc.
LLaVA-OneVision
Uniform 55.0 54.7 62.0 56.4 96.2
Cluster-based 55.6 55.2 62.4 57.3 97.2
Our DTM 55.9 56.3 62.7 57.3 98.0
LLaVA-Video
Uniform 55.4 54.6 62.9 59.8 91.5
Cluster-based 55.3 55.1 62.9 60.1 91.8
Our DTM 58.5 56.5 64.9 60.7 94.7
Results on LLaVA-Video. LLaVA-Video adopts a unique design by inserting newline tokens after each height-wise position in every frame, producing $64\times 13\times(13+1)$ tokens. In Table 2, for all methods, when the positional information of retained tokens is preserved, we also retain the associated newline tokens. For a fair comparison with VisionZip, we evaluate FastVID, where only one newline token is retained per frame. Notably, our FastVID consistently outperforms all baselines across various retention ratios.
[…]
4.3 Ablation Study
By default, we conduct ablation studies on LLaVA-OneVision at $r=10\%$. Further hyperparameter analysis is provided in the appendix.
Table 6: Ablation study on video segmentation.
Segmentation MVBench LongVideo- Bench MLVU VideoMME Avg. Acc.
Fixed-interval 55.1 53.6 61.7 55.4 95.3
Cluster-based 53.2 53.0 61.7 54.8 93.9
Our DySeg 55.9 56.3 62.7 57.3 98.0
Ablation study on DySeg. To compare different video segmentation methods, we present qualitative and quantitative results in Figure 3 and Table 6, respectively. In Table 6, Fixed-interval Segmentation with an interval of $4$ generates $32/4=8$ segments, while Cluster-based Segmentation generates $8$ clusters, ensuring consistency with the minimal segment number $c$ in our DySeg. Fixed-interval preserves temporal order between segments, whereas Cluster-based maintains intra-segment similarity. The results show that Fixed-interval outperforms Cluster-based by $1.4\%$, highlighting the importance of temporal structure preservation in Video LLM pruning. Additionally, high intra-segment similarity is also essential for effective pruning. Our proposed DySeg successfully integrates both advantages while enabling dynamic pruning, achieving a superior average performance of 98.0%.
Ablation study on STPrune. STPrune is proposed to prune tokens within each segment and consists of two key components: DTM and ATS. DTM employs density-based sampling to merge redundant tokens, preserving segment visual context. ATS leverages [CLS] attention scores to highlight important visual details. Table 5 presents an ablation study on $d$, which controls the pruning distribution between DTM and ATS. The results show that ATS alone (when $d=0.0$) outperforms DTM alone (when $d=1.0$). However, the best performance is achieved at $d=0.4$, demonstrating that a balanced integration of both modules is crucial for preserving essential video information. This integration boosts performance by 1.8% and 4.5% points over ATS and DTM alone, respectively.
Ablation study on DTM. In Table 5, we compare different token merging strategies for segment visual context. Uniform, used in VisionZip, selects anchor tokens in a content-agnostic manner, which limits its ability to distinguish semantically meaningful objects. Cluster-based discards essential positional information, resulting in inferior performance, particularly in LLaVA-Video. By leveraging Density-based Sampling and Anchor-centric Aggregation, our DTM achieves superior results, especially in LLaVA-Video, where DTM outperforms other methods by 2.9%.
Figure 5: Visualization of FastVID with $c=0,\tau=0.9,d=0.4,p=2$. We retain a total of 40 tokens, including 16 tokens in DTM (highlighted in red) and 24 in ATS (highlighted in blue). In DTM, patches that share the same inner and border color are merged.
4.4 Qualitative Results
Figure 5 visualizes the proposed DySeg and STPrune. In Figure 5(a), all frames are grouped into a single segment, with the 1st and 3rd frames selected as anchor frames, where DTM is applied. In Figure 5(b), the frames are divided into two segments, with the 1st, 3rd, and 4th frames selected as anchor frames. Blue tokens represent those selected by ATS due to high [CLS] attention. These tokens readily cluster together and do not always correspond to object regions, suggesting that while they reflect salient [CLS] information, they lack broader visual context. In contrast, DTM’s red tokens effectively aggregate visually similar content across frames, significantly reducing spatiotemporal redundancy while preserving visual context. Together, ATS and DTM offer complementary benefits for effective video compression. More qualitative results are provided in the appendix.
5 Conclusion
In this paper, we introduced FastVID, a novel inference-time pruning framework designed to accelerate Video LLMs by effectively reducing spatiotemporal redundancy. Through a comprehensive analysis of video tokens from both temporal context and visual context, FastVID dynamically partitions videos into temporally ordered segments and employs density-based token pruning within each segment. Extensive experiments across multiple Video LLMs and benchmarks demonstrate its generalization ability and effectiveness. Crucially, FastVID maintains high performance even under extreme compression rates, enabling practical deployment of fast Video LLMs.
Appendix
This appendix provides additional details and results to support our main paper:
  • Section A presents additional experimental settings, including reproduction details of the compared baselines and an estimation of the computational cost.
  • Section B presents additional experimental results, such as ablation studies on key hyperparameters, further qualitative results, and results on Qwen2-VL [37].
  • Section C discusses limitations of our approach and its potential broader impacts.
Appendix A Additional Experimental Settings
A.1 Reproduction Details of Compared Baselines
All experiments are conducted using LMMs-Eval [^2] https://github.com/EvolvingLMMs-Lab/lmms-eval, MIT License [44] for consistency. The performance of the vanilla versions of LLaVA-OneVision [^3] https://github.com/LLaVA-VL/LLaVA-NeXT, Apache License 2.0 [16], LLaVA-Video33footnotemark: 3 [47], and Qwen2-VL [^4] https://github.com/QwenLM/Qwen2.5-VL, Apache License 2.0 [37] differs slightly from their reported results, remaining within an acceptable margin of error. We reimplement all baseline methods using LMMs-Eval, following their official implementations:
  • FastV [^5] https://github.com/pkunlp-icler/FastV [5] (ECCV 2024). FastV performs token pruning at the $K$-th layer of the LLM using attention scores, with a filtering ratio $R$. We reimplement it with $K=2$, using $R\in\{75\%,80\%,85\%,90\%\}$ in Table 1 and $R\in\{75\%,90\%\}$ in Table 2.
  • VisionZip [^6] https://github.com/dvlab-research/VisionZip, Apache License 2.0 [42] (CVPR 2025). Visionzip performs pruning at the vision encoder’s output, which conflicts with pooling operations in Video LLMs and degrades performance. To address this, we implement VisionZip, which applies pruning after pooling. Following the original settings, each frame retains both dominant and contextual tokens in a $54\!:\!10$ ratio. We define $r$ as the proportion of tokens retained per frame. We set $r\in\{25\%,20\%,15\%,10\%\}$ in Table 1 and $r\in\{25\%,10\%\}$ in Table 2.
  • DyCoke [^7] https://github.com/KD-TAO/DyCoke, Apache License 2.0 [33] (CVPR 2025). DyCoke prunes in both prefilling and decoding stages. For fair comparison, we only evaluate its pre-filling stage. The pruning rate in the TTM module is set $K\in\{0.9,1.0\}$ in both Table 1 and Table 2.
  • PruneVID [^8] https://github.com/Visual-AI/PruneVid, CC BY-NC-SA 4.0 License [13] (ACL 2025). PruneVID includes both input-stage and intra-LLM pruning in the pre-filling phase, along with decoding pruning. For fair comparison, we evaluate two variants: PruneVID (input-stage pruning only) and PruneVID (full prefilling pruning). Following the original settings, we use a threshold $\tau=0.8$, temporal segment ratio $\gamma=0.25$, token selection ratio $\alpha=0.4$, and attention calculations use the 10th layer. The cluster ratio $\beta$ controls compression in the input-stage pruning. We use $\beta=11\%$ in Table 1 and Table 3.
A.2 Computational Cost Estimation
Following prior works [5, 39], we report the theoretical FLOPs of the LLM related to visual/video tokens. Specifically, LLaVA-OneVision [16], LLaVA-Video [47], and Qwen2-VL [37] are all built on Qwen2 [41], which employs grouped-query attention [2] and a three-layer FFN structure. The per-layer FLOPs of the LLM are computed as:
$2nD(h_{kv}d)+2nD^{2}+2n^{2}D+3nDD^{\prime}$ (6)
where $n$ is the number of video tokens, $D$ is the hidden state size, $D^{\prime}$ is the FFN intermediate size, $h_{kv}$ is the number of key/value heads, and $d$ is the head dimension.
Appendix B Additional Experimental Results
B.1 Ablation study on $c$ and $\tau$ in DySeg
We evaluate the effect of $c$ and $\tau$ in Eq. (2) of DySeg. Figure 6(a) shows results with varing $c$. In simple video scenarios with high transition similarity, $c$ regulates segmentation. When $c=1$, segmentation relies solely on $\tau$. $c=32$ divides the video into single-frame segments. Performance remains stable for $1\leq c\leq 16$, but the decline at $c=32$ suggests that excessive segmentation disregards temporal relationships. The optimal performance at $c=8$ indicates the benefit of a minimum cluster number. Figure 6(b) shows results with varing $\tau$. When $\tau$ is small, most transitions in $\mathbf{S}$ are selected by $c$. As $\tau$ increases, more transitions fall below the threshold. The best performance is achieved at $\tau=0.9$, effectively grouping redundant frames while separating non-redundant ones.
B.2 Ablation study on $p$ and $\beta$ in DTM
In Figure 7(a), we evaluate the effect of $p$ on anchor frame selection, from which anchor tokens are subsequently sampled. When $p=1$, anchor tokens are evenly distributed across all frames within a segment. As $p$ increases, fewer frames are selected as anchors, while more anchor tokens are allocated to each anchor frame. Given the high similarity between frames in a segment, a sufficient number of anchor tokens per anchor frame is essential to capture visual context. However, if $p$ is too large, all anchor tokens come from the first frame, limiting temporal information. We find that $p=4$ yields the best performance by effectively capturing spatiotemporal context. Figure 7(b) shows the effect of $\beta$ in Eq. (5). When $\beta=0.0$, matching tokens are averaged. $\beta=1.0$ discards all non-anchor tokens, removing segment visual context and leading to a performance drop. Notably, Anchor-centric Assignment (when $\beta=0.6$) yields optimal results, highlighting the importance of representative tokens and visual context.
Figure 6: Ablation study on $c$ and $\tau$ in DySeg. $c$ denotes the minimum number of segments in a video, whereas $\tau$ denotes the transition similarity threshold. Both parameters jointly control the segmentation granularity.
Figure 7: Ablation study on $p$ and $\beta$ in DTM. $p$ denotes the interval for anchor frame selection, whereas $\beta$ controls the merging weight for anchor tokens and their associated tokens in Eq. (5).
B.3 Additional Qualitative Results
Figures 8-10 show additional qualitative results on VideoMME. We use the following settings: $c=0,\tau=0.9,d=0.4,p=2$, retaining an average of 10 tokens per frame. Figure 8 presents static scenes, while Figures 9 and 10 present dynamic scenes. These visualizations demonstrate FastVID’s ability to adaptively segment videos across varying temporal dynamics. Within each segment, the merged region boundaries loosely align with object shapes, highlighting the effectiveness of the proposed DTM in preserving visual context.
Table 7: Comparison of state-of-the-art methods on Qwen2-VL [37]. A/B in # Tokens indicates that A tokens are provided as input to the LLM, and subsequently compressed to B tokens during the LLM forward pass. FastV performs pruning based on LLM attention scores, but its eager-attention implementation materializes the full attention matrix in memory, leading to OOM errors due to the large number of video tokens.
Method # Token TFLOPs VideoMME
Short Medium Long Overall
Vanilla 13447.1 100% 124.04 100% 74.1 60.4 54.3 63.0 100%
FastV [5]${}_{\text{ECCV'24}}$ 13447.1/3361.8 100%/25% 31.34 25.3% Out of Memory
VisionZip [42]${}_{\text{CVPR'25}}$ 3349.3 24.9% 24.11 19.4% 70.9 56.3 48.3 58.5 92.9%
PruneVID [13]${}_{\text{ACL'25}}$ 3460.6 25.7% 24.99 20.1% 66.7 54.0 48.1 56.3 89.4%
FastVID 3349.3 24.9% 24.11 19.4% 72.7 58.3 50.7 60.6 96.2%
B.4 Results on Qwen2-VL
Unlike LLaVA-OneVision [16] and LLaVA-Video [47], Qwen2-VL [37] employs a distinct architecture that samples 768 frames per video and processes them using 3D convolutions. It introduces M-RoPE, which decomposes rotary embeddings into temporal, height, and width components. The model dynamically adjusts the resolution of each frame. In Table 7, our FastVID outperforms other baselines on Qwen2-VL under similar model complexity. In particular, FastVID achieves a 2.4 accuracy improvement on the long subset. As shown in Tables 1, 2, and 7, our FastVID’s consistent superiority across three distinct Video LLM architectures further validates its strong generalizability and practical effectiveness.
Appendix C Additional Discussions
C.1 Limitations
FastVID achieves strong performance on LLaVA-OneVision [16] (32 frames/video), maintaining comparable accuracy with only 8.3% of the FLOPs. However, on LLaVA-Video [47] (64 frames/video) and Qwen2-VL [37] (768 frames/video), although FastVID consistently outperforms existing SoTA baselines, a noticeable accuracy drop occurs compared to the original model. We hypothesize that this is primarily due to the increased number of sampled frames, which leads to fewer tokens per frame and increases the semantic importance of each token. Consequently, aggressive pruning risks discarding critical information. In addition, although our query-agnostic token pruning effectively reduces spatiotemporal redundancy, it fails to emphasize query-relevant information sparsely distributed across extensively sampled frames. To overcome these challenges, future work may explore integrating text-guided keyframe selection to better support long-frame Video LLMs [37, 47].
C.2 Broader Impacts
FastVID accelerates inference for existing Video LLMs without modifying their parameters or training new models, thereby minimizing the risk of introducing new biases or unintended behaviors. However, FastVID inherits any potential negative societal impacts of the original models, such as representational bias or potential misuse. Despite this, FastVID maintains strong performance even under extreme token compression. This makes the use of Video LLMs more practical in environments with limited computational resources, promoting broader and more sustainable deployment.
Figure 8: Segment boundaries are marked by brown vertical lines. Tokens generated by DTM and ATS are highlighted in red and blue, respectively. Anchor frames, indicated by red boxes, are processed by DTM individually. In DTM, patches with matching inner and border colors are merged.
Figure 9: Segment boundaries are marked by brown vertical lines. Tokens generated by DTM and ATS are highlighted in red and blue, respectively. Anchor frames, indicated by red boxes, are processed by DTM individually. In DTM, patches with matching inner and border colors are merged.
Figure 10: Segment boundaries are marked by brown vertical lines. Tokens generated by DTM and ATS are highlighted in red and blue, respectively. Anchor frames, indicated by red boxes, are processed by DTM individually. In DTM, patches with matching inner and border colors are merged.